OpenStudy (anonymous):

Can someone plz help me!!!!!!!!!!! This question is do in 5 mins!!!

OpenStudy (anonymous):

What are the classifications of each system? Column A ---------------- 1. x+5y=-2 x+5y=4 2. y=3x+4 -2x+y=4 3. 5x+2y = -4 5x+2y=3 4. y= -x+3 2x+2y=6 --------------- Column B ---------------- A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

not C, hope this helps

OpenStudy (anonymous):

That didnt theres like four different questions here